OpenStudy (anonymous):

Multiplying a number by which of the following would result in a smaller number? a. 10^-1 b. 10^0 c. 10^1 d. 10^2 is it a?

Parth (parthkohli):

Actually, there are two answers to this :/

OpenStudy (anonymous):

AAAAAAAAAAAAAAA

OpenStudy (anonymous):

BBBBBBBBBBB IS ALSO

OpenStudy (anonymous):

oh ok thankyou!

OpenStudy (anonymous):

nooo not B :P

Parth (parthkohli):

\(10^2 = 100, 10^1 = 10, 10^0 = 1, 10^{-1} = 0.1\)

Parth (parthkohli):

Oops. Read the question differently.

OpenStudy (anonymous):

he says multiplying a number with one of these.. if u multiply any number by 10^0 u will get equal number not smaller

OpenStudy (anonymous):

A is right..

Parth (parthkohli):

If you multiply a number by 1, then you get the SAME number, not a number lesser.